OpenStudy (anonymous):

3n-40=1/2n+33 i got n = 1/(-73-sqrt(5335))

OpenStudy (anonymous):

how did u get that?

OpenStudy (anonymous):

neither of the n's are squared so how'd u get sqrt?

OpenStudy (anonymous):

it came out to 25 and i got that by trying to copy and paste it onto a caculator

OpenStudy (anonymous):

o i c what happened. is the equation 3n-40=(1/2)n+33 or 3n-40=1/(2n)+33

OpenStudy (anonymous):

?

OpenStudy (anonymous):

is it "a half of n" or "1 divided by 2n"?

OpenStudy (anonymous):

half

OpenStudy (anonymous):

\[3n-40 =1/2n+33 = (1+66n)/2n\] \[2n*(3n-40)=1+66n\] \[6n ^{2}-80n-66n-1=0\] \[6n ^{2}-146n-1 =0\] now solve for n using the quadratic formula

OpenStudy (anonymous):

3n-40=(1/2)n+33 2(3n-40)=2((1/2)n+33) 6n-80=n+66 6n-n-80=n-n+66 5n-80=66 5n-80+80-=66+80 5n=146 n=146/5 or 29+1/5

OpenStudy (anonymous):

thank you for steps both of you